OpenStudy (anonymous):

Can somebody please help me with this substitution problem??? 4x+7y=-11 -6x+y=51 Please explain:-)

Parth (parthkohli):

The easiest way you could use substitution is by solving for \(y\) in the second equation.\[y = 51 + 6x\]Now make the above substitution in the first equation.\[4x + 7(51 + 6x) = -11\]

OpenStudy (anonymous):

Ok so it should look like: 38x+357=-11

OpenStudy (helder_edwin):

\[ \large -11=4x+7(51+6x)=4x+357+42x=46x+357 \]

OpenStudy (anonymous):

Thanks a bunch!!!

OpenStudy (anonymous):

Sorry to bother you, but the answer that I got for x was -8...I keep getting a weird number for y, can you explain to me how you get it please?

OpenStudy (helder_edwin):

\[ \large 46x=-11-357=-368\Rightarrow x=-368/46=-8 \]

OpenStudy (helder_edwin):

\[ \large y=51+6x=51+6(-8)=51-48=3 \]